OpenStudy (emcheshhhh):

Part A: Sam rented a boat at $225 for 2 days. If he rents the same boat for 5 days, he has to pay a total rent of $480. Write an equation in the standard form to represent the total rent (y) that Sam has to pay for renting the boat for x days. (4 points) Part B: Write the equation obtained in Part A using function notation.(2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (wanell):

y = mx + C ----------------- When x=2, y=225. When x=5, y=480. ----------------- Therefore: 2m + C = 225 5m + C = 480 ------------------- 5*(2m+C)=5*225 2*(5m+C)=2*480 ------------------- Therefore: 10m + 5C = 1125 10m + 2C = 960 -------------- Therefore: 10m = 1125 - 5C 10m = 960 - 2C ------------- Therefore: 960 - 2C = 1125 - 5C 3C = 165 C = 55 --------------- y=mx + 55 --------------- 2m + C = 225 Therefore: 2m + 55 = 225 2m = 225 - 55 2m = 170 m = 85 ----------------- PART B: Which means that: y=85x + 55 ------------------- PART C & D: